Please Help me!!
Hi I desperately need help...the directions to this problem is to simplify
(xy^7)(x^4y)/(x^5y^2)
and i need help,, the directions are to add:
x-3/ x^2+7x+12
and i need help with solving this quadratic:
x^2-5x-1=0
if you can help me with explaining these topics in detail I would greatly appreciate your effort! Thank You!!

Answer by edjones(8007)

(xy^7)(x^4y)/(x^5y^2)
when multiplying add exponents; when dividing subtract them.
=(x^5y^11)/(x^5y^2)
=y^9
.
x-3/ x^2+7x+12
=(x-3)/(x+3)(x+4)
.
x^2-5x-1=0
x^2-5x =1
x^2-5x+6.25=6.25+1 Complete the square.
(x-2.5)^2=7.25
x-2.5=+-sqrt(7.25) sqrt of each side
x=2.5+sqrt(7.25), x=2.5-sqrt(7.25)
.
Ed